# Surface solar radiation: AI satellite retrieval can outperform Heliosat and generalizes well to other climate zones

### v1

@@ 5.2 Retrieval over Europe
Figure 2 illustrates differences and similarities across our five SSI retrieval models (Table 2) and Heliosat SARAH-3 on a randomly selected day and time. Overall, the five retrieval models mimic all cloud features seen in SARAH-3 at remarkable quality. When examining the retrievals closely, some differences become apparent. Firstly, the boundaries between clouds and clear-sky SSI tend to be somewhat sharper in SARAH-3 whereas the deep learning models exhibit moderate smoothing compared to SARAH-3. This may be attributed to the 16x16 pixel inputs which likely blend the SSI with neighboring pixel information. However, it is important to note that determining which level of smoothness is more accurate is challenging due to the absence of spatially extensive SSI measurements over large areas as ground truth.
@@ 5.3 Validating SSI retrieval models on ground stations
We validated our SSI retrieval models on multiple ground station networks across Europe and North Africa. To compare the models to actual SSI measurements rather than Heliosat SSI, we investigated root mean square errors (RMSE) and mean bias errors (MBE) with regard to the ground stations. The RMSE and MBE were calculated as averages over all matched ground station measurements from 2016 to 2022 including training set. Nominal start times HH:15 UTC and HH:45 UTC of the SEVIRI scan were excluded from the RMSE and MBE analysis to match the 30-minute frequency of the Heliosat SARAH-3 record. In Figures 5 and 6, the MBE and the difference in RMSE with regard to SARAH-3 were plotted by station. The IEA-PVPS dataset provides the largest coverage of the considered domain and is used for the analysis. The numeric values of MBE and RMSE for the IEA-PVPS dataset are given in Tables 3 and 4.
@@
Without fine-tuning, the emulator has a moderately higher RMSE (about 6-16 ) than Heliosat SARAH-3, as shown in Figure 6. Importantly, fine-tuning on any ground station network allows the emulator to outperform SARAH-3 SSI in terms of RMSE with regard to IEA PVPS ground stations. The initial stage of fine-tuning demonstrated performance improvements, as evidenced by reduced RMSEs when compared to the non-fine-tuned emulator across European stations (Table 3). The second and third stages of fine-tuning (on the MeteoSwiss and IEA PVPS stations) achieved a significant reduction in RMSE at nearly all sites of the four ground station networks, as shown in Figure 6. Remote stations such as in Tunisia, Algeria and Morocco exhibited RMSE reductions of 10 to 20 compared to SARAH-3. These improvements in RMSE and relatively stable MBE values suggest that our retrieval models generalize well to diverse geographical locations after fine-tuning on multiple ground station networks.
In most of central Europe, the MBE of model E (”Trained on DWD”) is low and its RMSE is close to that of the fine-tuned models, but for the stations to the south – the Spanish and four North African sites – it shows a bias of 50– 100 (Table 4) and worse accuracy than SARAH-3 at 30–47 higher RMSE (Table 3). The model that was fine-tuned on the DWD stations (”Finetuned DWD”, model B) did not show these large biases in the North African stations and improved in accuracy compared to the emulator.
@@ 5.3.1 Alpine validation
The emulator demonstrates a significant enhancement in performance across Alpine ground stations, as presented in Figures 5 and 6, particularly in regions with large RMSD relative to SARAH-3 (Figure 4(b)). Some of the ground stations, such as Gornergrat (GOR) at 3129 m a.s.l., are located at high altitudes on mountain ranges. Even though the emulator was trained solely on SARAH-3 SSI without any ground stations data, it achieves RMSE reductions of up to 50 W/m² compared to SARAH-3, indicating an improvement in retrieval accuracy.
Comparing SSI from Alpine stations such as Gornergrat to Heliosat SARAH-3 reveals a systematic underestimation of SSI by SARAH-3, as indicated by a pronounced off-diagonal frequency peak in Figure 12(a). This underestimation, which likely results from the misclassification of snow cover as clouds by Heliosat SARAH-3 [2], affects the majority of Alpine stations (not shown). Notably, the emulator does not replicate this pattern, even without fine-tuning (Figure 12(b)), which suggests that the spectral signatures of snow cover differ sufficiently from those of clouds to allow for improved distinction.
Two anomalies in the retrieval of Alpine SSI are evident in Figure 5, involving models B (”Finetuned on DWD”) and E (”Trained on DWD”). Both models were trained on DWD stations and display significant biases at certain Alpine stations. Model B, which was fine-tuned on DWD data, exhibits a bias exceeding 150 W/m² across the entire Alps. The scatter plot for model B (”Finetuned on DWD”) at GOR (Figure 12(B)) shows that this bias stems from an overestimation at low SSI values. Conversely, model E (”Trained on DWD”) exhibits a substantial negative bias (¡-150 W/m²) at some Alpine stations, including GOR, as reflected by the underestimation in the scatter plot (Figure 12(E)). Similar patterns are observed at other high-altitude stations. One possible explanation for the underestimation with model E is that the spectral signature of Alpine surfaces under clear-sky conditions differs markedly from the surfaces predominant in the DWD training set.
@@ 5.3.2 Representativeness error
Meteosat SEVIRI Level-1.5 data are provided at a typical grid size of 3x3 km, so the SSI estimates represent instantaneous values of the mean SSI in a 3x3 km area. In contrast, SSI ground observations provide point measurements that are averaged over periods of 1-10 minutes and have limited spatial representativeness. Huang et al. [46] evaluated the representative errors for an SSI product on a grid scale of 5x5 km. Seventeen pyranometer stations within a 5x5 km area at a location in Northern China measured subgrid scale deviations. At a 10-minute averaging of ground observations, they found the average representativeness error of grid-scale products to be with an average RMSE of for the SSI product. With increasing cloud cover fraction, this error increases to around with an RMSE of . Clouds increase the representativeness error due to subgrid inhomogeneity, but also solar retrieval products do not take into account 3D radiative transfer effects within clouds [46]. While representativeness errors may, in principle, affect our validation, we have demonstrated above that fine-tuning our retrieval models of instantaneous SSI on 10-minute averaged ground observations still increases the performance by up to . This shows that, despite potential representativeness errors, our deep-learning retrieval models learn systematic spatial SSI patterns to provide more accurate representative SSI retrieval.
@@ 5.4 Cloudy versus clear-sky conditions
The performance of our SSI retrieval models is discussed for different cloudiness conditions in the following. SSI retrieval errors depend on the level of cloudiness present. Quesada-Ruiz et al. [17] showed that the RMSE of Heliosat-2 SSI with regard to ground stations depends on the CSI. They presented a retrieval method that achieved the largest reductions in RMSE in overcast conditions, where the Heliosat retrievals had the highest RMSE. Likewise, Verbois et al. [20] accomplished the largest improvements in RMSE for samples with CSI ¡ 0.6. We investigate the performance of our five retrieval models in clear-sky and cloudy conditions by separating the ground station measurements into CSI classes. The clear-sky index was calculated from measured SSI at the ground stations and using the Ineichen-Perez clear-sky surface radiation model [47]. Figure 7 shows SSI error statistics for different levels of cloudiness as expressed by the CSI. Heliosat SARAH-3 underestimates SSI by around -25 in clear-sky index larger than 1 and overestimates SSI by around 27–90 for CSI ¡ 0.8. The large systematic biases in SARAH-3 and other radiative-transfer-based SSI retrievals (positive for CSI¡0.8 and negative for CSI1) [48] may partially be related to fractional clouds (Jörg Trentmann, personal communication). Clouds that partially cover the grid cell are measured at a point location while the satellite observes an average. The clear-sky index defined on the basis of ground observations and not over the entire grid cell can result in these biases. The emulator and the fine-tuned models mimick the negative and positive biases of Heliosat SARAH-3 (Figure 7). While the emulator generates somewhat larger residuals, fine-tuning on ground stations effectively decreased the residual spread for CSI¡0.8 compared to SARAH-3.
@@
Interestingly, solely training on ground stations without any Heliosat emulation (model E ”Trained on DWD”) generates the smallest biases in cloudy conditions (CSI¡0.8). The model trained only on ground stations (model E) also features the best generalizability to locations that are remote from its training set sites. Therefore, in cloudy conditions, the model trained solely on DWD stations provides SSI estimates with the smallest biases and RMSEs even at out-of-domain locations compared to the other models and with large performance improvements over Heliosat (Figure 7). The superior SSI retrieval performance of model E is exemplified in Figure 13 for the Moroccan IEA-PVPS station. Even though the model was trained only on ground stations in Germany without any training data from any other regions, it shows the best SSI retrieval performance in cloudy conditions even in countries such as Morocco that are thousands of kilometers from the training set domain and feature very different climates and surface albedos. The model shows the best performance with an improvement of 40 in RMSE with respect to SARAH-3. The other North African and Spanish stations exhibit the same effect. The respective performance in terms of MBE and RMSE on the IEA-PVPS dataset for cloudy CSI ¡ 0.8 and clear-sky CSI 0.8 samples is depicted in Tables 3 and 4. Models trained on ground stations of DWD and MeteoSwiss show in general a worse RMSE on the North-African and Spanish stations in clear-sky conditions than SARAH-3, while in cloudy conditions the performance improves compared to SARAH-3. In cloudy conditions, the model trained solely on DWD stations (model E, ”Trained on DWD”) has the lowest RMSE and lowest absolute MBE for all IEA-PVPS stations (Figure 7), demonstrating the model’s generalizability to out-of-domain locations.
@@
Our results indicate that the generalizability of a data-driven SSI retrieval model is not only related to the model training data or method, but also depends on the level of cloudiness present in the location for which SSI is retrieved with the data-driven model. This observation can provide additional explanation for the insufficient SSI retrieval generalizability seen by Verbois et al. [20] and Yang et al. [24].
@@ 5.5 Importance of SEVIRI channels for estimating SSI
Our study is the first, to our knowledge, to quantify how relevant the individual channels of a geostationary imager are for accurate SSI retrieval. As the channels relate to different wavelength bands, they differ in their relative importance for capturing atmospheric composition aspects and surface properties [35]. For example, the visible VIS006 and VIS008 are more relevant for cloud opacity due to reflectivity whereas WV062 and WV073 mainly sense tropospheric water vapor. We quantified the relevance of the individual SEVIRI channels in an observations-based, data-driven manner based on the permutation feature importance of the predictor variables of each of our five SSI retrieval models. Permutation feature importance methods assess the importance of each predictor variable (feature) in a machine learning model based on how much the predictor variable contributes to the model’s performance [49]. Permutation feature importance quantifies how the model performance changes when the values of a predictor variable are randomly permuted while keeping the values of the other predictor variables unchanged. If the random shuffling leads to a significant drop in model accuracy, the predictor variable whose values were shuffled is important for the model’s performance. After permuting a predictor variable over the entire validation set, we calculated the model performance in terms of the mean absolute error (MAE) and compared it to the model performance without permutation. The 15x15 grids represent a single feature per channel in the feature importance.
@@
The permutation feature importances are shown in Figure 8. Shuffling a predictor variable that has no effect on the retrieval model’s performance, such as WV062, leaves the MAE practically unchanged. Shuffling features essential to the retrieval model – such as the VIS006 channel and the SZA – causes large performance decreases of around 100 and 200 in MAE, respectively. The predictor variables latitude, longitude, azimuth, DEM and day of the year were left out from Figure 8 as permutation feature importance showed they had no significant effect on the model performance. The emulator effectively disregards the VIS008 channel. However, the other deep learning retrieval models retrieve information and gain performance from the VIS008 channel. The emulator shows the lowest feature importance for all channels but attributes the largest importance to the SZA, compared to the other models, which suggests it does not yet make optimal use of the information contained in the SEVIRI channels.
@@
Figures 10 and 11 illustrate predictor importance for clear-sky and cloudy conditions. As expected, our SSI retrieval models rely mostly on SZA in clear-sky conditions, because the SZA includes the majority of information to estimate SSI without cloud attenuation. In cloudy conditions, the visible, near-infrared and infrared channels gain importance. For model E (”Trained on DWD”), the IR087 channel plays a significant role in the clear-sky conditions compared to the other retrieval models. IR087, and a combination of IR108 and IR120, are important for distinguishing ice clouds from water clouds [35] and may be helpful in distinguishing snow cover from clouds in clear-sky conditions. We expect that a data-driven evaluation of channel importance can contribute to further improvement of radiative-transfer based models.
@@ 6 Conclusions
Our study introduced the first machine-learning-based satellite retrieval for instantaneous surface solar irradiance and demonstrated its capability to provide accurate and generalizable SSI estimates across Europe. Our retrieval models were trained on multiple years of Meteosat SEVIRI observations in eleven visible and infrared channels. Heliosat SSI estimates and SSI from ground station networks were used as target variables in the machine learning model development. Instantaneous SSI is not averaged in time and primarily of interest in short-term forecasting, such as intra-hour solar forecasts for applications in solar energy production and building control. Our study is the first to emulate a radiative-transfer based retrieval model. We introduced a deep learning emulator of a Heliosat SSI retrieval (SARAH-3) and showed that the emulator is an accurate surrogate of thereof. By including SSI measurements from ground stations in the model training, our SSI retrieval model outperformed Heliosat in terms of accuracy and biases in cloudy conditions (CSI ¡ 0.8). When trained on ground station SSI, our retrieval models generalized well to regions with other climates and surface albedos. Our results indicate that the generalizability of a data-driven SSI retrieval model is not only related to the model training data or training method, but also depends on the amount of cloudiness present in the location at which SSI is retrieved with the data-driven model. We found that, in cloudy conditions, a model trained only on ground stations can estimate SSI accurately even in locations with different surface characteristics, far away from the training test domain. Our SSI retrieval model trained on ground stations provides out-of-domain generalization and achieves large performance gains compared to Heliosat. Moreover, we found emulation is helpful but not necessary to achieve good generalizability. In fact, for cloudy conditions (CSI¡0.8), better generalizability (lowest bias and RMSE) was achieved by direct model training on ground stations.
@@
We also showed that the SSI of Heliosat SARAH-3 exhibits large biases in mountain regions, and that training and fine-tuning our retrieval models on ground station SSI strongly reduces these large biases, outperforming Heliosat. Further, our study quantified the relevance of the Meteosat channels and other predictor variables, such as solar zenith angle and time of year, for the model accuracy in different cloud conditions. Permutation feature importance showed that in cloudy conditions, multiple near-infrared and infrared channels enhance performance. Our results can facilitate the development of more accurate radiative-transfer-based SSI retrieval models. Future research should investigate additional predictor variables and their impact on the retrieval quality, notably to improve the performance in clear-sky conditions. Fore example, including aerosol-related observations or estimates of clear-sky irradiance can hold the potential to enhance the DL retrieval model’s accuracy also in clear-sky conditions (CS¿0.8).

### v2

@@ 5.2 Retrieval over Europe
Figure 2 illustrates differences and similarities across our five SSI retrieval models (Table 4.4) and Heliosat SARAH-3 on a randomly selected day and time. Overall, the five retrieval models mimic all cloud features seen in SARAH-3 at remarkable quality. When examining the retrievals closely, some differences become apparent. Firstly, the boundaries between clouds and clear-sky SSI tend to be somewhat sharper in SARAH-3 whereas the deep learning models exhibit moderate smoothing compared to SARAH-3. This may be attributed to the 1616 pixel inputs which likely blend the SSI with neighboring pixel information. However, it is important to note that determining which level of smoothness is more accurate is challenging due to the absence of spatially extensive SSI measurements over large areas as ground truth.
@@ 5.3 Validating SSI retrieval models on ground stations
We validated our SSI retrieval models on multiple ground station networks across Europe and North Africa. To compare the models to actual SSI measurements rather than Heliosat SSI, we investigated root-mean-square errors (RMSE) and mean bias errors (MBE) with regard to the ground stations. The RMSE and MBE were calculated as averages over all matched ground station measurements from 2016 to 2022 including training set. Nominal start times HH:15 UTC and HH:45 UTC of the SEVIRI scan were excluded from the RMSE and MBE analysis to match the 30-minute frequency of the Heliosat SARAH-3 record. In Figures 5 and 6, the MBE and the difference in RMSE with regard to SARAH-3 were plotted by station. The IEA-PVPS dataset provides the largest coverage of the considered domain and is used for the analysis. The numeric values of MBE and RMSE for the IEA-PVPS dataset are given in Tables B and B.
@@
Without fine-tuning, the emulator has a moderately higher RMSE (about 6-16 ) than Heliosat SARAH-3, as shown in Figure 6. Importantly, fine-tuning on any ground station network allows the emulator to outperform SARAH-3 SSI in terms of RMSE with regard to IEA PVPS ground stations. The initial stage of fine-tuning demonstrated performance improvements, as evidenced by reduced RMSEs when compared to the non-fine-tuned emulator across European stations (Table B). The second and third stages of fine-tuning (on the MeteoSwiss and IEA PVPS stations) achieved a significant reduction in RMSE at nearly all sites of the four ground station networks, as shown in Figure 6. Remote stations such as in Tunisia, Algeria and Morocco exhibited RMSE reductions of 10–20 compared to SARAH-3. These improvements in RMSE and relatively stable MBE values suggest that our retrieval models generalize well to diverse geographical locations after fine-tuning on multiple ground station networks.
In most of Central Europe, the MBE of model E (’Trained on DWD’) is low and its RMSE is close to that of the fine-tuned models, but for the stations to the south – the Spanish and four North African sites – it shows a bias of 50–100 (Table B) and worse accuracy than SARAH-3 at 30–47 higher RMSE (Table B). The model that was fine-tuned on the DWD stations (’Finetuned DWD’, model B) did not show these large biases in the North African stations and improved in accuracy compared to the emulator.
@@ 5.3.1 Alpine validation
The emulator demonstrates a significant enhancement in performance across Alpine ground stations, as presented in Figures 5 and 6, particularly in regions with large RMSD relative to SARAH-3 (Figure 4). Some of the ground stations, such as Gornergrat (GOR) at 3129 m a.s.l., are located at high altitudes on mountain ranges. Even though the emulator was trained solely on SARAH-3 SSI without any ground stations data, it achieves RMSE reductions of up to 50 W/m² compared to SARAH-3, indicating an improvement in retrieval accuracy.
Comparing SSI from Alpine stations such as Gornergrat to Heliosat SARAH-3 reveals a systematic underestimation of SSI by SARAH-3, as indicated by a pronounced off-diagonal frequency peak in Figure 12(a). This underestimation, which likely results from the misclassification of snow cover as clouds by Heliosat SARAH-3 (Carpentieri et al. 2023), affects the majority of Alpine stations (not shown). Notably, the emulator does not replicate this pattern, even without fine-tuning (Figure 12(b)), which suggests that the spectral signatures of snow cover differ sufficiently from those of clouds to allow for improved distinction.
Two anomalies in the retrieval of Alpine SSI are evident in Figure 5, involving models B (’Finetuned on DWD’) and E (’Trained on DWD’). Both models were trained on DWD stations and display significant biases at certain Alpine stations. Model B, which was fine-tuned on DWD data, exhibits a bias exceeding 150 W/m² across the entire Alps. The scatter plot for model B (’Finetuned on DWD’) at GOR (Figure 12(B)) shows that this bias stems from an overestimation at low SSI values. Conversely, model E (’Trained on DWD’) exhibits a substantial negative bias (-150 W/m²) at some Alpine stations, including GOR, as reflected by the underestimation in the scatter plot (Figure 12(E)). Similar patterns are observed at other high-altitude stations. One possible explanation for the underestimation with model E is that the spectral signature of Alpine surfaces under clear-sky conditions differs markedly from the surfaces predominant in the DWD training set.
@@ 5.3.2 Representativeness error
Meteosat SEVIRI Level-1.5 data are provided at a typical grid size of 33 km, so the SSI estimates represent instantaneous values of the mean SSI in a 33 km area. In contrast, SSI ground observations provide point measurements that are averaged over periods of 1-10 minutes and have limited spatial representativeness. Huang et al. (2016) evaluated the representative errors for an SSI product on a grid scale of 55 km. Seventeen pyranometer stations within a 55 km area at a location in Northern China measured subgrid scale deviations. At a 10-minute averaging of ground observations, they found the average representativeness error of grid-scale products to be with an average RMSE of for the SSI product. With increasing cloud cover fraction, this error increases to around with an RMSE of . Clouds increase the representativeness error due to subgrid inhomogeneity, but also solar retrieval products do not take into account 3D radiative transfer effects within clouds (Huang et al. 2016). While representativeness errors may, in principle, affect our validation, we have demonstrated above that fine-tuning our retrieval models of instantaneous SSI on 10-minute averaged ground observations still increases the performance by up to . This shows that, despite potential representativeness errors, our deep-learning retrieval models learn systematic spatial SSI patterns to provide more accurate representative SSI retrieval.
@@ 5.4 Cloudy versus clear-sky conditions
The performance of our SSI retrieval models is discussed for different cloudiness conditions in the following. SSI retrieval errors depend on the level of cloudiness present. Quesada-Ruiz et al. (2015) showed that the RMSE of Heliosat-2 SSI with regard to ground stations depends on the CSI. They presented a retrieval method that achieved the largest reductions in RMSE in overcast conditions, where the Heliosat retrievals had the highest RMSE. Likewise, Verbois et al. (2023a) accomplished the largest improvements in RMSE for samples with CSI 0.6. We investigate the performance of our five retrieval models in clear-sky and cloudy conditions by separating the ground station measurements into CSI classes. The clear-sky index was calculated from measured SSI at the ground stations and using the Ineichen-Perez clear-sky surface radiation model (Perez et al. 1990). Figure 7 shows SSI error statistics for different levels of cloudiness as expressed by the CSI. Heliosat SARAH-3 underestimates SSI by around -25 in clear-sky index larger than 1 and overestimates SSI by around 27–90 for CSI 0.8. The large systematic biases in SARAH-3 and other radiative-transfer-based SSI retrievals (positive for CSI 0.8 and negative for CSI 1) (Ineichen et al. 2009) may partially be related to fractional clouds (Jörg Trentmann, personal communication). Clouds that partially cover the grid cell are measured at a point location while the satellite observes an average. The clear-sky index defined on the basis of ground observations and not over the entire grid cell can result in these biases. The emulator and the fine-tuned models mimic the negative and positive biases of Heliosat SARAH-3 (Figure 7). While the emulator generates somewhat larger residuals, fine-tuning on ground stations effectively decreased the residual spread for CSI 0.8 compared to SARAH-3.
@@
Interestingly, solely training on ground stations without any Heliosat emulation (model E ’Trained on DWD’) generates the smallest biases in cloudy conditions (CSI 0.8). The model trained only on ground stations (model E) also features the best generalizability to locations that are remote from its training set sites. Therefore, in cloudy conditions, the model trained solely on DWD stations provides SSI estimates with the smallest biases and RMSEs even at out-of-domain locations compared to the other models and with large performance improvements over Heliosat (Figure 7). The superior SSI retrieval performance of model E is exemplified in Figure 13 for the Moroccan IEA-PVPS station. Even though the model was trained only on ground stations in Germany without any training data from any other regions, it shows the best SSI retrieval performance in cloudy conditions even in countries such as Morocco that are thousands of kilometers from the training set domain and feature very different climates and surface albedos. The model shows the best performance with an improvement of 40 in RMSE with respect to SARAH-3. The other North African and Spanish stations exhibit the same effect. The respective performance in terms of MBE and RMSE on the IEA-PVPS dataset for cloudy CSI 0.8 and clear-sky CSI 0.8 samples is depicted in Tables B and B. Models trained on ground stations of DWD and MeteoSwiss show in general a worse RMSE on the North-African and Spanish stations in clear-sky conditions than SARAH-3, while in cloudy conditions the performance improves compared to SARAH-3. In cloudy conditions, the model trained solely on DWD stations (model E, ’Trained on DWD’) has the lowest RMSE and lowest absolute MBE for all IEA-PVPS stations (Figure 7), demonstrating the model’s generalizability to out-of-domain locations.
@@
Our results indicate that the generalizability of a data-driven SSI retrieval model is not only related to the model training data or method, but also depends on the level of cloudiness present in the location for which SSI is retrieved with the data-driven model. This observation can provide additional explanation for the insufficient SSI retrieval generalizability seen by Verbois et al. (2023a) and Yang et al. (2022b).
@@ 5.5 Importance of SEVIRI channels for estimating SSI
Shi et al. (2023) showed first the importance of individual channels and attributes to the retrieval of SSI. As the channels relate to different wavelength bands, they differ in their relative importance for capturing atmospheric composition aspects and surface properties (Schmetz et al. 2002). For example, the visible VIS006 and VIS008 are more relevant for cloud opacity due to reflectivity whereas WV062 and WV073 mainly sense tropospheric water vapor. Here we continue the trend of explaining the importance of features but for neural network-type models. We quantified the relevance of the individual SEVIRI channels in an observations-based, data-driven manner based on the permutation feature importance of the predictor variables of each of our five SSI retrieval models. Permutation feature importance methods assess the importance of each predictor variable (feature) in a machine learning model based on how much the predictor variable contributes to the model’s performance (Fisher et al. 2019). Permutation feature importance quantifies how the model performance changes when the values of a predictor variable are randomly permuted while keeping the values of the other predictor variables unchanged. If the random shuffling leads to a significant drop in model accuracy, the predictor variable whose values were shuffled is important for the model’s performance. After permuting a predictor variable over the entire validation set, we calculated the model performance in terms of the mean absolute error (MAE) and compared it to the model performance without permutation. The 1515 grids represent a single feature per channel in the feature importance.
@@
The permutation feature importance is shown in Figure 8 based on the IEA-PVPS dataset. Shuffling a predictor variable that does not affect the retrieval model’s performance, such as WV062, leaves the MAE practically unchanged. Shuffling features essential to the retrieval model – such as the VIS006 channel and the SZA – cause large performance decreases of around 100 and 200 in MAE, respectively. The predictor variables latitude, longitude, azimuth, DEM and day of the year were left out from Figure 8 as permutation feature importance showed they had no significant effect on the model performance. In Shi et al. (2023) the solar azimuth angle shows significant importance. The difference might be that the surrounding pixel information given to the CNN might make the solar azimuth angle redundant. The emulator effectively disregards the VIS008 channel. However, the other deep learning retrieval models retrieve information and gain performance from the VIS008 channel. The emulator shows the lowest feature importance for all channels but attributes the largest importance to the SZA, compared to the other models, which suggests it does not yet make optimal use of the information contained in the SEVIRI channels.
@@
Figures 10 and 11 illustrate predictor importance for clear-sky and cloudy conditions. As expected, our SSI retrieval models rely mostly on SZA in clear-sky conditions, because the SZA includes the majority of information to estimate SSI without cloud attenuation. In cloudy conditions, the visible, near-infrared and infrared channels gain importance. For model E (’Trained on DWD’), the IR087 channel plays a significant role in the clear-sky conditions compared to the other retrieval models. IR087, and a combination of IR108 and IR120, are important for distinguishing ice clouds from water clouds (Schmetz et al. 2002) and may be helpful in distinguishing snow cover from clouds in clear-sky conditions. We expect that a data-driven evaluation of channel importance can contribute to further improvement of radiative-transfer based models.
@@ 6 Conclusions
Our study introduced a machine-learning-based satellite retrieval for instantaneous surface solar irradiance and demonstrated its capability to provide accurate and generalizable SSI estimates across Europe. Our retrieval models were trained on multiple years of Meteosat SEVIRI observations in eleven visible and infrared channels. Heliosat SSI estimates and SSI from ground station networks were used as target variables in the machine learning model development. Instantaneous SSI is not averaged in time and primarily of interest in short-term forecasting, such as intra-hour solar forecasts for applications in solar energy production and building control. Our study is the first to emulate a radiative-transfer based retrieval model. We introduced a deep learning emulator of a Heliosat SSI retrieval (SARAH-3) and showed that the emulator is an accurate surrogate of thereof. By including SSI measurements from ground stations in the model training, our SSI retrieval model outperformed Heliosat in terms of accuracy and biases in cloudy conditions (CSI 0.8). When trained on ground station SSI, our retrieval models generalized well to regions with other climates and surface albedos. Our results indicate that the generalizability of a data-driven SSI retrieval model is not only related to the model training data or training method, but also depends on the amount of cloudiness present in the location at which SSI is retrieved with the data-driven model. We found that, in cloudy conditions, a model trained only on ground stations can estimate SSI accurately even in locations with different surface characteristics, far away from the training test domain. Our SSI retrieval model trained on ground stations provides out-of-domain generalization and achieves large performance gains compared to Heliosat. Moreover, we found emulation is helpful but not necessary to achieve good generalizability. In fact, for cloudy conditions (CSI 0.8), better generalizability (the lowest bias and RMSE) was achieved by direct model training on ground stations. While transfer learning did not increase performance for cloudy conditions, it did help with spatial generalization in clear-sky conditions where the model trained solely on ground stations predicted SSI with large biases.
@@
We also showed that the SSI of Heliosat SARAH-3 exhibits large biases in mountain regions, and that training and fine-tuning our retrieval models on ground station SSI strongly reduces these large biases, outperforming Heliosat. Further, our study quantified the relevance of the Meteosat channels and other predictor variables, such as solar zenith angle and time of year, for the model accuracy in different cloud conditions. Permutation feature importance showed that in cloudy conditions, multiple near-infrared and infrared channels enhance performance. Our results can facilitate the development of more accurate radiative-transfer-based SSI retrieval models. Future research should investigate additional predictor variables and their impact on the retrieval quality, notably to improve the performance in clear-sky conditions or to replace the clear-sky retrieval with a physical model to create a hybrid model.
Our study shows it is possible to emulate a RTM-based retrieval model with DL. Finetuning this base model on ground stations will lead to a better generalizable model with high performance but there are limitations and difficulties to the methodology. Finetuning requires quite a bit of trial and error and advanced training methods from the field of transfer learning to not overfit the new type of dataset. Unforeseen biases in particular climates, such as mountainous regions of the domain, can still occur using a finetuned model. The general accuracy gain of up to 20% compared to SARAH-3 in cloudy samples should prove the usefulness of DL retrieval since those are most important for short-term solar forecasts. We expect our contribution will facilitate faster, better and more generalizable SSI retrieval with better knowledge of the pitfalls DL retrieval brings.